OpenStudy (anonymous):

PLEASE HELP! (Waiting for an hour..._) Find an equation of the tangent line to the curve at the given point?? \(\ \Large \frac{x^2}{16}-\frac{y^2}{9} =1\) \(\ \Large \text{The point is: } (-5, \frac{9}{4}) \). PLEASE HELP!

OpenStudy (anonymous):

what math are you in?

OpenStudy (anonymous):

AP Calculus BC

OpenStudy (anonymous):

what's "BC"

OpenStudy (anonymous):

It's just a level of calculus.... I need to find the derivative of this equation...

OpenStudy (anonymous):

solve for y and take the derivative?

OpenStudy (anonymous):

This section involves implicit differentiation

OpenStudy (anonymous):

ohhh

OpenStudy (anonymous):

did you take the derivative?

OpenStudy (anonymous):

No I'm still stuck on this problem!!

OpenStudy (anonymous):

I don't know how to with those fractions

OpenStudy (anonymous):

the derivative of \(\frac{x^2}{16}\) is \(\frac{x}{8}\)

OpenStudy (anonymous):

Really? Wouldn't the 16 become a zero?

OpenStudy (anonymous):

and the derivative with respect to \(x\) of \(-\frac{y^2}{9}\) is \[-\frac{2y}{9}y'\]

OpenStudy (anonymous):

no it is a constant, think \[\frac{x^2}{16}=\frac{1}{16}x^2\]

OpenStudy (anonymous):

Am I just using power rule here? I don't need to use quotient rule (that's what I was thinking...?)

OpenStudy (anonymous):

@satellite73 I've been getting mixed answers from people. When do I know when I take the derivative of y to have yy' versus just y'???